consider all samples of size 4 from this population: 1,5,6,8,10. calculate the standard deviation of the sampling distribution of the sample means


1
Expert's answer
2022-03-30T01:47:46-0400


mean, "\mu=6.0"

Standard deviation "\sigma=\sqrt{x^2f(x)-(6)^2}"

"=\sqrt{36.575-36}"

"=\sqrt{0.575}"

"=0.7583"